# SLED: A Speculative LLM Decoding Framework for Efficient Edge Serving

### v4

@@ IV-A Whole System Token Generation Rate (WSTGR)
IV-C Cost Efficiency and System Throughput
Figure 6: Pareto front showing optimal trade-offs between energy consumption per token and latency, highlighting the efficiency of SLED.
The cost efficiency of token generation in edge computing scenarios is a critical factor and has been considered in various edge inference system designs [30, 31]. In this paper, we compare the proposed SLED system, a centralized serving system, and an all-edge decoding system in terms of both cost efficiency and WSTGR. To systematically analyze the trade-off between cost and performance, we construct Pareto front visualizations, which highlight the non-dominated configurations that achieve the best balance between monetary cost and system throughput.
In our experiments, the cost and throughput metrics for edge inference scenarios were carefully computed based on a comprehensive capital expenditure (CAPEX) and operational expenditure (OPEX) model[32]. Specifically, we adopt the widely recognized CPU-hour cost model described by Walker[32] and the edge-compute modeling approach proposed by Eriksson[33]. The CAPEX component was determined by amortizing the purchase price of each edge device (Raspberry Pi 5 priced at \$80 [34]) over a three-year lifetime, assuming an average device utilization rate of 70%. The OPEX component included electrical consumption calculated from experimentally measured average power draw (8 W for Pi 5) [35] and industrial electricity rates (0.083 \$/kWh) [36]. Combining these costs, we obtained a unified hourly expense for each device, subsequently normalized by the experimentally measured token generation rates (tokens per second), as shown in Eq. 2. The resulting metric, expressed clearly as dollars per one thousand generated tokens (\$/1K tokens), enabled direct and transparent comparison across different experimental configurations and devices.
$\text{Cost}\,=\frac{1000}{3600\,R}\,\Bigl{(}\frac{P_{\text{device}}}{3\times 8760\times 0.70}\;+\;\frac{P_{\text{avg}}}{1000}\times 0.083\Bigr{)}$ (2)
@@ IV-C Cost Efficiency and System Throughput
V Conclusion and Future Work
This position paper presented the SLED, a novel distributed decoding framework designed for LLM deployment at the edge. Our extensive evaluation demonstrated that SLED significantly improves system throughput, capacity, and cost efficiency compared to traditional centralized approaches. The integration of speculative local drafting and centralized verification establishes a balance of computational workload, making SLED particularly suitable for bringing LLMs towards the edge of the network. It’s highlighted in this position paper that the SLED is more than a decoding enhancement—it opens the door to a more foundational and elastic approach to resource-aware LLM serving at the edge.
Future research directions include exploring the adaptive queue and batching strategy on the edge server for latency-sensitive tasks and better server utilization. Additionally, enhancing the adaptive capabilities of SLED for dynamic environments, such as extending SLED’s applicability to multi-modal scenarios, will be another interesting topic to focus on. Lastly, network conditions and resource-aware verification strategy could further broaden its practical impact in complex edge computing landscapes.
VI Acknowledgment
This material is based on work supported by the National Science Foundation under Grants No. 2315851 and 2106634

### v5

@@ 4.1. Whole System Token Generation Rate (WSTGR)
4.3. Cost Efficiency and System Throughput
Figure 6. Pareto front showing optimal trade-offs between energy consumption per token and latency, highlighting the efficiency of SLED.:
The cost efficiency of token generation in edge computing scenarios is a critical factor and has been considered in various edge inference system designs (Husom2025SustainableLLM; Jang2025EdgeFirst). In this paper, we compare the proposed SLED system, a centralized serving system, and an all-edge decoding system in terms of both cost efficiency and WSTGR. To systematically analyze the trade-off between cost and performance, we construct Pareto front visualizations, which highlight the non-dominated configurations that achieve the best balance between monetary cost and system throughput.
In our experiments, the cost and throughput metrics for edge inference scenarios were carefully computed based on a comprehensive capital expenditure (CAPEX) and operational expenditure (OPEX) model(walker2009cpu). Specifically, we adopt the widely recognized CPU-hour cost model described by Walker(walker2009cpu) and the edge-compute modeling approach proposed by Eriksson(eriksson2020cost). The CAPEX component was determined by amortizing the purchase price of each edge device (Raspberry Pi 5 priced at \$80 (upton2023pi5)) over a three-year lifetime, assuming an average device utilization rate of 70%. The OPEX component included electrical consumption calculated from experimentally measured average power draw (8 W for Pi 5) (odonnell2023review) and industrial electricity rates (0.083 \$/kWh) (eia2025electric). Combining these costs, we obtained a unified hourly expense for each device, subsequently normalized by the experimentally measured token generation rates (tokens per second), as shown in Eq. 3. The resulting metric, expressed clearly as dollars per one thousand generated tokens (\$/1K tokens), enabled direct and transparent comparison across different experimental configurations and devices.
(3) $\text{Cost}\,=\frac{1000}{3600\,R}\,\Bigl(\frac{P_{\text{device}}}{3\times 8760\times 0.70}\;+\;\frac{P_{\text{avg}}}{1000}\times 0.083\Bigr)$
@@ 4.3. Cost Efficiency and System Throughput
4.4. Impacts of Network Connectivity
We built a SLED prototype to quantify how network loss affects edge-side throughput. The edge runs on a Raspberry Pi 5B hosting meta-llama/Llama-3.2-1B-Instruct for drafting, while the server uses A100 GPU running meta-llama/Meta-Llama-3.1-70B-Instruct for batched verification. We instrumented the system to sweep packet-loss ratios from 0% to 100% under a fixed RTT budget and measured the committed token throughput at the edge. As shown in Fig.7, across commonly used settings in practice ($\leq$5–10% loss), SLED’s edge throughput degraded only slightly (typically ¡2–3% versus the no-loss baseline), confirming that the proposed retry-then-fallback policy effectively masks moderate loss. Even under complete network unavailability (responses never arrive), the edge maintained $\geq$5.24 tokens/s, comfortably above typical human reading rates, ensuring responsive user-perceived progress. These results indicate that, in both engineering practice and research deployments, SLED preserves acceptable edge throughput under realistic packet-loss regimes, while providing graceful performance under extreme conditions.
Figure 7. The impact of packet loss on edge device-side throughput.:
Additionally, we evaluated the response quality of the SLED under varying packet loss rates in the prototype system. By sweeping the packet loss ratio from 0% to 100%, we assessed performance on GSM8K (Cobbe2021GSM8K), a curated benchmark of linguistically diverse, grade-school math word problems designed to evaluate multi-step quantitative reasoning in language models. As shown in Fig.8, the response quality remains consistent with that of the target model when the verification packet loss rate is below 10%, which covers the most common real-world conditions. The quality gradually declines to the draft model level as the packet loss rate increases from 10% to 100%, indicating a loss of connection to the verification server. Based on the proposed retry-then-fallback policy, persistent packet loss triggers a fallback to the local draft model for token generation, leading to a degradation in output quality. These results demonstrate that under typical connectivity conditions, SLED can effectively serve edge devices without compromising accuracy.
Figure 8. The impact of packet loss on the quality of generated response.:
@@ 4.4. Impacts of Network Connectivity
5. Conclusion and Future Work
This position paper presented the SLED, a novel distributed decoding framework designed for LLM deployment at the edge. Our extensive evaluation demonstrated that SLED significantly improves system throughput, capacity, and cost efficiency compared to traditional centralized approaches. The integration of speculative local drafting and centralized verification establishes a balance of computational workload, making SLED particularly suitable for bringing LLMs towards the edge of the network. The position paper highlights that the SLED is more than a decoding enhancement—it opens the door to a more foundational and elastic approach to resource-aware LLM serving at the edge.
In future work, we will explore optimizing key–value caching strategies, potentially leveraging recent advances in PagedAttention, to further reduce server-side verification cost. Additionally, enhancing the adaptive capabilities of SLED for dynamic environments, such as extending SLED’s applicability to multi-modal scenarios, will be another interesting topic to focus on. Lastly, network conditions and resource-aware verification strategy could further broaden its practical impact in complex edge computing landscapes.
Acknowledgements.
This material is based on work supported by the National Science Foundation under Grants No. 2315851 and 2106634.